# Compress then Serve: Serving Thousands of LoRA Adapters with Little Overhead

### v2

@@ 7 Discussion
This study introduces approaches to LoRA compression, addressing significant challenges facing foundation models and large language models. Our contributions include theoretical formulations, empirical validation, and practical implementations that enhance the understanding and application of LLMs in scalable environments.
The implications of our findings are manifold. Our theoretical guarantees for reconstruction error not only increase confidence in the use of compressed models but also lay a groundwork for future explorations in this area. Demonstrating that our compression techniques can preserve up to 100% of the original LoRAs’ performance highlights the effectiveness of our methods. Furthermore, integrating LoRA compression into state-of-the-art LLM serving systems demonstrates the potential for resource optimization, with throughput for thousands of LoRAs nearing that of a single LoRA.
The promising results of our study suggest several future research directions. First, further compression may be possible via quantization. Our joint-diagonalization compression and quantization are independent axes of approaching the problem and exploring a combined solution can be fruitful. Second, when scaling to hundreds of thousands of LoRAs, joint compression, while effective, will not be sufficient to fit all LoRAs onto the GPU, thus requiring a procedure to schedule the requests. Our clustering variant offers opportunities to develop an efficient scheduling mechanism that takes into account the cluster assignments of LoRAs corresponding to the incoming requests.
In conclusion, our research significantly advances the deployment of LLMs by providing robust, scalable, and efficient compression solutions. The ability of compressed LoRAs to maintain high performance while facilitating substantial resource savings opens new avenues for the broader application and adoption of LLMs across various industries. We encourage the community to build upon our findings and the shared LoRAs to further explore and enhance the utility of these technologies.
@@ A.1 Alternating Methods
[…]
@@ A.3 Clustering algorithm
Initialization: We run joint diagonalization with a single $U,V$ then perform k-means with $|C_{j}|$ clusters on the space of $\Sigma_{i}$’s. This gives us our first clusters and we can use random initialization $U_{j},V_{j}$ for each cluster but the $\Sigma_{i}$ can be maintained as initialization.
Step 1: Using the alternating JD algorithms from earlier in this section, we optimize the problem $\min_{U_{j},V_{j},\Sigma_{i}}\sum_{i\in C_{j}}||B_{i}A_{i}-U_{j}\Sigma_{i}V_{j}^{\top}||^{2}_{F}$ for each $j$ independently.
Step 2: New cluster assignment for $i$ : $\min_{j}\min_{\Sigma_{i}}||B_{i}A_{i}-U_{j}\Sigma_{i}V_{j}^{\top}||^{2}_{F}$. If any assignment changes we go to Step 1, else we have converged.

### v3

@@ 7 Discussion
This study introduces approaches to LoRA compression, addressing significant challenges emerging as customization of foundation models such as LLMs and diffusion models becomes increasingly popular. Our contributions include theoretical formulations, empirical validation, and practical implementations that enhance the understanding and application of LLMs in scalable environments.
Our findings have several implications. Our theoretical bounds on reconstruction error not only increase confidence in the use of compressed models but also lay a groundwork for future explorations. Demonstrating that our compression techniques can preserve up to 100% of the original LoRAs’ performance highlights the effectiveness of our methods. Furthermore, integrating LoRA compression into state-of-the-art LLM serving systems demonstrates potential for resource optimization, with throughput for thousands of LoRAs nearing that of a single LoRA.
Our promising results suggest several future research directions. First, further compression may be possible via quantization, since joint-diagonalization and quantization are independent compression strategies. Second, when scaling to hundreds of thousands of LoRAs, joint compression, while effective, will be insufficient to fit all LoRAs onto the GPU, thus requiring a procedure to schedule the requests. Clustering offers opportunities for efficient scheduling that incorporates the cluster assignments of LoRAs corresponding to the incoming requests.
Privacy presents another research direction, particularly regarding the possibility of information leakage during joint compression. As a preliminary study, Appendix H.2 investigates whether a base model with an adapter $A$ for task $T_{A}$, after being jointly compressed alongside an adapter $B$ for task $T_{B}$, inadvertently improves on $T_{B}$. Such an outcome would indicate that adapter $A$ acquired information from adapter $B$. Our ablation study shows no performance gains on $T_{B}$, suggesting that the compressed adapter $A$ remains independent and does not leak—or gain—information from adapter $B$. A more detailed investigation of the privacy properties of joint compression is an interesting next step.
In conclusion, our research advances LLM deployment by providing robust, scalable, and efficient compression. The ability of compressed LoRAs to maintain high performance while saving resources opens avenues for the broad application and adoption of LLMs across various industries. We encourage the community to build upon our findings and shared LoRAs to further enhance these technologies.
Impact Statement
This paper presents work whose goal is to advance machine learning. There are no societal consequences of our work that we feel must be specifically highlighted here.
@@ A.1 Alternating Methods
$=\textrm{const.}-2\sum_{i}\mathrm{tr}(B_{i}A_{i}V\Sigma_{i}^{\top}U^{\top})+\sum_{i}\|U\Sigma_{i}V^{\top}\|_{\mathrm{Fro}}^{2}.$ (5)
@@ A.3 Clustering algorithm
Initialization: We run joint diagonalization with a single $U,V$ then perform k-means with $k$ clusters on the space of $\Sigma_{i}$’s. This gives us our first clusters and we can use random initialization $U_{j},V_{j}$ for each cluster but the $\Sigma_{i}$ can be maintained as initialization.
[…]
@@ Appendix C Training LoRAs
In Table 1, we compare metrics between base model and LoRA finetuning.
Table 1: Comparison of metrics before and after LoRA training across 1000 tasks.
Metric Base Model LoRA
Loss $4.14\pm 3.07$ $0.56\pm 0.58$
Exact Match $1.81\pm 6.56$ $51.38\pm 40.90$
Rouge-1 $21.70\pm 19.22$ $68.88\pm 29.73$
Rouge-L $20.62\pm 18.21$ $67.80\pm 30.15$